OpenStudy (anonymous):

Compute your average velocity in the following ase: (a) You walk 78.8 m at a speed of 1.91 m/s and then run 78.8 m at a speed of 3.57 m/s along a straight track.

OpenStudy (anonymous):

you should find the time of walking in two stages and add them . v=d/t 1.91=78.8/t1 t1=78.8/1.91=41.2 s v=d/t 3.57=78.8/t2 t2=78.8/3.57=22.07s t=t1+t2=41.2+22.07=63.27 s